A person wants to resolve two thin poles standing near each other from the distance of 2km.The minimum separation between poles should be nearly (theta=1 arc min )
1) 30 cm
2)40cm
3)60 cm
4)120 cm
1
Expert's answer
2021-09-09T10:50:12-0400

"\theta= 1\text{arc min}"

"1\text{arc min}= \frac{1}{60\degree}=\frac{1}{60}*\frac{\pi}{180}\approx0.0003\text{ rad}"

"r = 2km = 2000m"

"\theta=\frac{l}{r}"

"l =\theta r=0.0003*2000= 0.6m = 60cm"


"\text{Answer: }3)60 cm"